OpenStudy (anonymous):

write the expression in radical notation (16^1/2)^-3

OpenStudy (anonymous):

\[(16^{\frac{1}{2}})^{-3}=\frac{1}{16^\frac{3}{2}}=\frac{1}{\sqrt[2]{16^3}}\]

OpenStudy (jack1):

(16^1/2)^-3 \[16^{1/2} = \sqrt{16} \] (x)^-3 \[x^{-3} = \frac {1 }{x^3}\]

OpenStudy (anonymous):

Just to point out, \[\sqrt{16^3}=(\sqrt{16})^3\]